OpenStudy (anonymous):

find the point-slope form then change to ax + by = c using m=2 and (2,4)

OpenStudy (michele_laino):

hint: we can write: \[y - 4 = 2\left( {x - 2} \right)\]

OpenStudy (michele_laino):

so, next step is: \[y - 2 = 2x - 4\]

OpenStudy (anonymous):

ok so y-2=2x-4 is the point slope form then? now i need to change it to ax+by= c form how do i do that

OpenStudy (anonymous):

subtract \(y\) from both sides, then add \(4\) to both sides

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ 4}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= 2 \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)